OpenStudy (anonymous):

Solve: cos(x)(cosx+1)=0

OpenStudy (anonymous):

is it cos(x+1)?

OpenStudy (anonymous):

or (cos x)+1

OpenStudy (anonymous):

its (cosx+1)

OpenStudy (anonymous):

like...i...said...

OpenStudy (anonymous):

either cos(x) is 0 or cos(x+1) is 0

OpenStudy (anonymous):

Set each to 0 and solve \[\cos(x) = 0\] \[\cos(x) + 1 = 0 or \cos(x) = -1\] cos(x) is 0 when x is pi/2 3pi/2, 5pi/2 ect so the answer would be \[x = \frac{\pi}{2} +\pi*n\] and cos(x) is -1 when x is pi, 3pi, 5pi,ect so the answer to that would be \[x = \pi + 2n*\pi\]

OpenStudy (anonymous):

oh, i should mention, n is a integer value

OpenStudy (anonymous):

The avalaible answers are A. x = PI/3+-2PIn, x = 3PI/4+-2PIn B. x = Pi/2+-2PIn, x = 3PI/2P-2PIn, x =Pi+-2PIn C. x =+-PIn, x = Pi/2+-2PIn D. x = Pi/2+-PIn, x = 3Pi/2+-2Pin

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Wrong again but hey..tbh im not really surprized...everyone has given me wrong answers for the last 6 questions...why should i expect dif

OpenStudy (anonymous):

wait hang on..

OpenStudy (anonymous):

too late moving on thnx anyways tho

OpenStudy (anonymous):

It was b ._.

OpenStudy (anonymous):

ohh, i see what they did in B. They did it weird.

OpenStudy (anonymous):

yeah well..im starting to get pissed ...one question wrong occasionally im fine with....but the last 6 ..well now 7 have gotten wrong

OpenStudy (anonymous):

in.a.row. .....

OpenStudy (anonymous):

Life is tough